# Boosted Neural Decoders: Achieving Extreme Reliability of LDPC Codes for 6G Networks

### v1

@@ IV-A Evaluation Environment
Training is carried out using TensorFlow and the Adam optimizer [47], on an NVIDIA GeForce RTX A5000 GPU and an AMD EPYC 7763 CPU. For fast simulation during the sampling process and FER evaluation, the C programming language is employed. To guarantee the uniqueness of each random experiment, a long-cycle noise generator, specifically the Mersenne Twister algorithm [48], is utilized. The FERs are evaluated using the Monte Carlo method, with experiments repeated until a minimum of $100$ UC vectors are identified.
@@ IV-B Result for WiMAX and WiFi LDPC Codes
We compare the performance with the MS decoding, WMS decoding, and other algorithm-based decoding approaches targeting the error floor [11, 12]. The TAMS decoder [11] with the parameter set $\alpha=0.8,\tau=3.0$ shows a similar performance with the WMS and MS decoders in the waterfall and error floor regions, respectively. We also compare with the state-of-the-art post processing scheme [12]. The simulation results are directly referenced. While the post processing scheme [12] shows comparable or slightly inferior performance to the boosted NMS decoder, it suffers from high decoding complexity and latency due to its large number of iterations ¯$\overline{\ell}=150$. On the contrary, the boosted NMS decoder exhibits a barely noticeable error floor down to FER of $10^{-7}$ with ¯$\overline{\ell}=50$ iterations. In the comparison between ¯$\overline{\ell}=20$ and ¯$\overline{\ell}=50$ settings for the boosted NMS decoder, it is evident that the post decoder successfully mitigates the error floor.
@@
Fig. 9(b) shows the FER performances for the WiFi LDPC code with $R=5/6,n=648$. In this experiment, the boosted NMS decoder is compared with conventional NMS decoders [9, 20]. We exclude neural decoders that rely on different decoding algorithms or that require enumerating trapping sets. For the NMS decoders in [9, 20], the base decoding stage is the same as that of the boosted NMS decoder, while the post decoding stage is trained by their respective training methods. The training method in [9] involves the use of received vectors from the waterfall region (${\rm E}_{\rm b}/{\rm N}_{0}$ $2$-$4$ dB) as training samples, training all weights all at once without a training schedule. For the training method in [20], received vectors from the ${\rm E}_{\rm b}/{\rm N}_{0}$ points where the MS decoder achieves BER of $10^{-3}$ are used as training samples, and the iter-by-iter training schedule is employed. Both methods use the full weight set to achieve their best performance. The other hyper-parameters are aligned with those in our proposed method. As shown in Fig. 9(b), the conventional NMS decoders [9, 20] perform well in the waterfall region (${\rm E}_{\rm b}/{\rm N}_{0}$ $2$-$4$ dB), but start to exhibit the error floor beyond $4$ dB. In contrast, the boosted NMS decoder shows excellent performance in both the waterfall and error floor regions.
Table VII offers a comparison of the decoding complexity for the WiMAX LDPC code. The boosted NMS decoder has additional multiplications for the weighting operation by $(E+N)z$ compared to the MS decoder and by $Nz$ compared to the WMS and TAMS [11] decoders. The number of other operations remains the same as that of the others. The overall complexity is evaluated with the assumption that the comparison operation $C$ is twice as complex as the addition $A$ and multiplication $M$ [49]. The boosted NMS decoder requires additional memory for storing $2\ell_{1}+3\ell_{2}$ weights, whose number is less than the number of weights for the conventional NMS decoders [9, 20]. Since the post-processing scheme in [12] does not use weighting, the complexity per iteration is lower than the boosted NMS scheme. However, the total complexity exceeds that of the proposed method by more than double due to the higher required number of iterations. Furthermore, additional complexity is required for the error path detector [12].
@@
TABLE VII: Complexity comparison for the WiMAX LDPC code of $(N,M,E,z,\alpha)=(24,6,88,24,15.6)$, where $\alpha$ is the comparison count for each CN (i.e., $\alpha=d_{c}+\lceil{\rm log}d_{c}\rceil-2$ for the CN degree $d_{c}$ [43])
@@ IV-C Results for 5G LDPC Codes
Figure 10: FER performances of (a): $(R=0.5,n=576)$ and $(R=0.5,n=714)$, (b): $(R=0.33,n=816)$ and $(R=0.75,n=372)$ 5G LDPC codes.
@@
In Fig. 10, we evaluate the performance of four 5G LDPC codes with the following specifications: $(R=0.5,n=576),(R=0.5,n=714),(R=0.33,n=816),(R=0.75,n=372)$. For 5G LDPC codes that exhibit weak protection in the parity parts, we specifically evaluate the FER for information parts only. To train the boosted NMS decoder that achieves the xURLLC FER target of $10^{-9}$, we utilize the methods of transfer learning and data augmentation. The total of $50$ iterations is divided into $20$ for the base stage, and $15$ each for the first and second post stages. Since the first post decoder with only $15$ iterations is able to correct most of UC vectors from the base decoder, we do not allocate additional iterations nor employ the block-wise training schedule. Instead, the second post decoder is used to correct UC vectors from the base and first post decoders. In training the second post decoder, we use a more flexible weight set $\{\overline{w}^{(\ell)}_{v_{p}},w^{(\ell)}_{c_{p}},{\hat{w}}^{(\ell)}_{c_{p}}\}_{\ell}$ to further improve performance.
@@
The base decoder is trained in the waterfall region as usual. For the $(R=0.5,n=576)$ code, the first post decoder is trained using $10{,}000$ UC vectors without applying transfer learning. We refer to the trained weights as the baseline weights. These baseline weights are then used in transfer learning for training the second post decoder for the $(R=0.5,n=576)$ code, as well as the first and second post decoders for the other codes. For data augmentation, $100$ UC vectors are collected at the onset of the error floor region and augmented into $1{,}000$ UC vectors.
@@
Fig. 10 demonstrates that the boosted NMS decoder achieves the xURLLC target for all evaluated codes without encountering any severe error floor. For example, with the $(R=0.5,n=714)$ code, while the MS decoder achieves the target at $6.5$ dB, the boosted NMS decoder achieves it at $4.8$ dB, showing an improvement of $1.7$ dB. In addition, it shows superior performance in the waterfall region as well. For other codes, the boosted NMS decoder also achieves the target FER at lower ${\rm E}_{\rm b}/{\rm N}_{0}$ values. Owing to the accelerated sampling speed by transfer learning and data augmentation, it is possible to train the boosted NMS decoder to achieve the extremely low error rate of $10^{-9}$ across various codes.
@@ IV-D Comparison with Augmented Neural Decoders
TABLE VIII: Comparison with [9, 23, 39] in terms of waterfall performance of short codes (The results are measured by the negative natural logarithm of BER. Higher is better.)
@@
In Table VIII, we compare with augmented neural decoders [23, 39] for short length codes including polar, BCH, MacKay LDPC codes in terms of waterfall performance. To ensure a fair comparison with other works, we employ the BP decoding algorithm, which results in the boosted NBP decoder, and utilize the soft-BER loss function for BER performance optimization while making use of a general weight set $\{\overline{w}_{v}^{(\ell)},w_{c\rightarrow v}^{(\ell)},\hat{w}_{c\rightarrow v}^{(\ell)}\}_{\ell}$.
Table VIII shows a notable performance enhancement over the original NBP decoder [9], which employs the standard NBP architecture. The improvement becomes more significant at higher SNR, where the boosted NBP decoder even outperforms the decoder [39] that incorporates hyper-networks into the model architecture. When compared to the ECCT [23] using the transformer architecture, our method shows inferior performance for high density codes (Polar, BCH codes), yet it achieves fairly comparable performance for LDPC codes. Notably, the boosted NBP decoder with ¯$\overline{\ell}=100$ outperforms the ECCT with $N=10$ for the MacKay LDPC code, where $N$ refers to the number of layers in the transformer architecture.
In Fig. 11, the boosted NBP decoder is compared with the ECCT [23] as well as the BP-RNN decoder [22] for a short LDPC code of length $128$, which is provided in [22]. The BP-RNN decoder employs noisy codeword vectors generated with the knowledge of absorbing sets for its training samples. The performance data is directly referenced from [22]. For a fair comparison, the boosted NMS decoder is configured similarly to the parallel BP-RNN decoder [22] where the total of ¯$\overline{\ell}=250=10\times 25$ iterations is segmented into $10$ parallel stages, each with $25$ iterations. Thus, it can be performed with low latency decoding in parallel. This configuration includes $1$ base decoder and $9$ post decoders in the context of boosting learning. The result shows that the boosted NBP decoder with ¯$\overline{\ell}=10\times 25$ surpasses both the BP-RNN decoder [22] and ECCT [23]. This underlines the effectiveness of the proposed training methodology not just in the error floor region of long LDPC codes, but also in the waterfall region of short LDPC codes.
Even though the comparison in this subsection considers augmented neural networks as competitors, we want to note that our proposed training methodology is adaptable regardless of decoding architectures. This flexibility presents a promising way for future research, particularly in terms of integrating our training techniques with advanced architecture like hyper-networks or the transformer architecture. Such integration could potentially enhance the decoder performance for various code types.

### v2

@@ IV-A Evaluation Environment
Training is carried out using the Adam optimizer [56] with TensorFlow, on an NVIDIA GeForce RTX A5000 GPU and an AMD EPYC 7763 CPU. For fast simulation during the sampling process and FER evaluation, the C programming language is employed. To guarantee the uniqueness of each random experiment, a long-cycle noise generator, specifically the Mersenne Twister algorithm [57], is utilized. The FER performances are evaluated using the Monte Carlo method, with experiments repeated until a minimum of $100$ UC vectors are identified.
@@ IV-B Result for WiMAX and WiFi LDPC Codes
We compare the performance with the MS decoding, WMS decoding, offset MS (OMS) decoding, and other algorithm-based decoding approaches targeting the error floor [11, 12]. The TAMS decoder [11] with the parameter set $\alpha=0.8,\tau=3.0$ shows a similar performance with the WMS and MS decoders in the waterfall and error floor regions, respectively. We also compare with the state-of-the-art post processing scheme [12]. Simulation results are directly referenced. While the post processing scheme [12] shows comparable or slightly inferior performance to the boosted NMS decoder, it suffers from high decoding complexity and latency due to its large number of iterations ¯$\overline{\ell}=150$. On the contrary, the boosted NMS decoder exhibits a barely noticeable error floor down to FER of $10^{-7}$ with ¯$\overline{\ell}=50$ iterations. In the comparison between ¯$\overline{\ell}=20$ and ¯$\overline{\ell}=50$ settings for the boosted NMS decoder, it is evident that the post decoder successfully mitigates the error floor.
@@
Fig. 9(b) shows the FER performances for the WiFi LDPC code with $R=5/6,n=648$. In this experiment, the boosted NMS decoder is compared with conventional NMS decoders [9, 20]. We exclude neural decoders that rely on different decoding algorithms or require the enumeration of trapping sets. For the NMS decoders in [9, 20], the base decoding stage is the same as that of the boosted NMS decoder, while the post decoding stage is trained by their respective training methods. The training method in [9] involves the use of received vectors from the waterfall region (${\rm E}_{\rm b}/{\rm N}_{0}$ $2$-$4$ dB) as training samples, training all weights all at once without a training schedule. For the training method in [20], received vectors from the ${\rm E}_{\rm b}/{\rm N}_{0}$ points where the MS decoder achieves BER of $10^{-3}$ are used as training samples, and the iter-by-iter training schedule is employed. Both methods use the full weight set to achieve their best performance. The other hyper-parameters are aligned with those in our proposed method. As shown in Fig. 9(b), the conventional NMS decoders [9, 20] perform well in the waterfall region (${\rm E}_{\rm b}/{\rm N}_{0}$ $2$-$4$ dB), but start to exhibit the error floor beyond $4$ dB. In contrast, the boosted NMS decoder shows excellent performance in both the waterfall and error floor regions.
@@
TABLE VIII: Complexity comparison for the WiMAX LDPC code of $(N,M,E,z,\alpha)=(24,6,88,24,15.6)$, where $\alpha$ is the comparison count for each CN (i.e., $\alpha=d_{c}+\lceil{\rm log}d_{c}\rceil-2$ for the CN degree $d_{c}$ [52])
@@
Table VIII offers a comparison of the decoding complexity for the WiMAX LDPC code. The boosted NMS decoder has additional multiplications for the weighting operation, requiring $(E+N)z$ more than the MS decoder and $Nz$ more than the WMS and TAMS [11] decoders. The number of other operations remains unchanged. The overall complexity is evaluated with the assumption that the comparison operation $C$ is twice as complex as the addition $A$ and multiplication $M$ [58]. The boosted NMS decoder requires additional memory for storing $2\ell_{1}+3\ell_{2}$ weights, whose number is less than the number of weights for the conventional NMS decoders [9, 20]. Since the post-processing scheme in [12] does not use weighting, the complexity per iteration is lower than that of the boosted NMS scheme. However, the total complexity exceeds that of the proposed method by more than double due to the higher required number of iterations. Furthermore, additional complexity is introduced by the error path detector [12].
@@ IV-C Results for 5G LDPC Codes
Figure 10: FER performances of (a): $(512,256)$ and $(1024,512)$ 5G LDPC codes, (b): $(786,256)$ and $(352,256)$ 5G LDPC codes, (c): $(512,256)$ 5G LDPC code with various message representations and (c): $(2112,1536)$ 5G LDPC code.
@@
In Figs. 10(a) and (b), we evaluate the performance of four $(n,k)$ 5G LDPC codes with the following specifications: $(512,256)$, $(1024,512)$, $(786,256)$, $(352,256)$, whose code rates are $0.5,0.5,0.33,0.73$, respectively. For 5G LDPC codes that exhibit weak protection in the parity parts, we specifically evaluate FER for the information parts only. To train the boosted NMS decoder that achieves the xURLLC FER target of $10^{-9}$, we utilize the methods of transfer learning and data augmentation. The total $50$ iterations are divided into $20$ for the base stage, and $15$ each for the first and second post stages. Since the first post decoder with only $15$ iterations is able to correct most UC vectors from the base decoder, we do not allocate additional iterations nor employ the block-wise training schedule. Instead, the second post decoder is used to correct UC vectors from the base and first post decoders. In training the second post decoder, we use a more flexible weight set $\{\overline{w}^{(\ell)}_{v_{p}},w^{(\ell)}_{c_{p}},{\hat{w}}^{(\ell)}_{c_{p}}\}_{\ell}$ to further improve performance.
@@
The base decoder is trained in the waterfall region as usual. For the $(512,256)$ code, the first post decoder is trained using $10{,}000$ UC vectors without applying transfer learning. We denote the trained weights as baseline weights. These baseline weights are then used in transfer learning for training the second post decoder for the $(512,256)$ code, as well as the first and second post decoders for the other codes. For data augmentation, $100$ UC vectors are collected at the onset of the error floor region and augmented into $1{,}000$ UC vectors.
@@
Figs. 10(a) and (b) demonstrate that the boosted NMS decoder achieves the xURLLC target across all evaluated codes without encountering any severe error floor. For example, with the $(512,256)$ code, the boosted NMS decoder reaches the target FER $10^{-9}$ at $5$ dB, while the WMS decoder only achieves FER of $10^{-5}$ at the same $5$ dB, implying $10^{4}$ gain by the boosted NMS decoder. In addition, it shows superior performance in the waterfall region as well. For other codes, the boosted NMS decoder also achieves the target FER at lower ${\rm E}_{\rm b}/{\rm N}_{0}$ values. Owing to the accelerated sampling speed by transfer learning and data augmentation, it is possible to train the boosted NMS decoder to achieve the extremely low error rate of $10^{-9}$ across various codes.
One reason the WMS decoder exhibits a severe error floor starting from FER of $10^{-4}$, as seen in Figs. 10(a) and (b), is the quantization scheme. To investigate the impact of quantization, Fig. 10(c) presents the performance of different message representations. Up until now, we have assumed a quantization scheme with a step size of $\Delta=0.5$ to achieve finer message representation, which is effective in the waterfall region. However, this representation imposes a maximum LLR value of $7.5$, leading to poor error floor performance. Increasing the step size to $1.0$ allows the maximum LLR value to rise to $15.0$, resulting in a more gradual error floor starting from $10^{-5}$, though at the cost of degraded waterfall performance, as shown in Fig. 10(c). If quantization is eliminated and floating-point message representations are used instead, both the waterfall and error floor performances improve significantly, but this is impractical for hardware implementation. In contrast to this trade-off for the WMS decoder, the boosted NMS decoder shows a balanced performance in both the waterfall and error floor regions for all message representations. Notably, the quantized boosted NMS decoder (blue and green curves) nearly aligns with the boosted NMS decoder using floating-point implementation (red curve) and outperforms the BP decoder at low error rates, implying that the negative effects of quantization are mitigated by boosting learning. This is of significant practical importance, as quantization is essential for cost-effective hardware implementation.
Fig. 10(d) shows the performances of the $(2112,1536)$ long 5G LDPC code, where the BP decoder operating with floating-point message representation exhibits an error floor below $10^{-5}$. Although 5G LDPC codes are designed to avoid an error floor above FER of $10^{-4}$ with the BP decoder [59], they do not guarantee the absence of an error floor below this point. When the proposed techniques are applied to the BP algorithm, resulting in the boosted NBP decoder, the error floor is almost entirely mitigated. Similarly, the boosted NMS decoder also achieves improved error floor performance compared to the WMS decoder. This result demonstrates that the proposed methods can be broadly applied to various decoding algorithms.
@@ IV-D Comparison with Augmented Neural Decoders
TABLE IX: Comparison with [9, 23, 48] in terms of waterfall performance of short codes (The results are measured by the negative natural logarithm of BER. Higher is better.)
@@
In Table IX, we compare with augmented neural decoders [23, 48] for short length codes including polar, BCH, MacKay LDPC codes in terms of waterfall performance. To ensure a fair comparison with other works, we employ the boosted NBP decoder and utilize the soft-BER loss function for BER performance optimization while making use of a general weight set $\{\overline{w}_{v}^{(\ell)},w_{c\rightarrow v}^{(\ell)},\hat{w}_{c\rightarrow v}^{(\ell)}\}_{\ell}$.
Table IX shows a notable performance enhancement over the original NBP decoder [9], which employs the standard NBP architecture. The improvement becomes more significant at higher SNR levels, where the boosted NBP decoder even outperforms the decoder [48] that incorporates hyper-networks into the model architecture. When compared to the ECCT [23] using the transformer architecture, our method shows inferior performance for high density codes (Polar, BCH codes), but achieves fairly comparable performance for LDPC codes. Notably, the boosted NBP decoder with ¯$\overline{\ell}=100$ outperforms the ECCT with $N=10$ for the MacKay LDPC code, where $N$ refers to the number of layers in the transformer architecture.
In Fig. 11, the boosted NBP decoder is compared with the ECCT [23] as well as the BP-RNN decoder [22] for a short LDPC code of length $128$, which is provided in [22]. The BP-RNN decoder employs noisy codeword vectors generated with the knowledge of absorbing sets for its training samples. The performance data is taken directly from [22]. For a fair comparison, the boosted NMS decoder is configured similarly to the parallel BP-RNN decoder [22] where the total of ¯$\overline{\ell}=250=10\times 25$ iterations is segmented into $10$ parallel stages, each with $25$ iterations. Thus, it can be performed with low latency decoding in parallel. This configuration includes $1$ base decoder and $9$ post decoders in the context of boosting learning. The result shows that the boosted NBP decoder with ¯$\overline{\ell}=10\times 25$ surpasses both the BP-RNN decoder [22] and ECCT [23]. This underlines the effectiveness of the proposed training methodology not just in the error floor region of long LDPC codes, but also in the waterfall region of short LDPC codes.
Even though the comparison in this subsection considers augmented neural networks as competitors, we want to note that our proposed training methodology is adaptable regardless of decoding architectures. This flexibility opens up promising avenues for future research, particularly in terms of integrating our training techniques with advanced architectures such as hyper-networks and the transformer architectures. Such integration could potentially enhance the decoder performance for various code types.